OpenStudy (anonymous):

Need 1 help Plz help me........

OpenStudy (anonymous):

what is it

OpenStudy (anonymous):

If \[1^{2}+2^{2}+3^{2}+.......+512^{2}=m\]\[how is 2^{2}+4^{2}+6^{2}........+1024^{2}=4m\] Plz explain

OpenStudy (anonymous):

you can use the formula for the sum of squares i believe do you know it?

OpenStudy (anonymous):

no bro @satellite73

OpenStudy (anonymous):

i.e. \[\sum_{k=1}^n k^2=\frac{n(n+1)(2n+1)}{6}\]

terenzreignz (terenzreignz):

Or straight into the fire :P \[n=2^2 + 4^2 + 6^2 +...1024^2\\=(2\cdot 1)^2 +(2\cdot2)^2 + (2\cdot 3)^2+...(2\cdot512)^2\\=4\cdot1^2 + 4\cdot2^2+4\cdot3^2 + ...4\cdot512^2\] And you get the idea, hopefully :D

OpenStudy (anonymous):

so \[\sum_{k=1}^{512}=\frac{512(513)(1025)}{6}=m\]

OpenStudy (anonymous):

actually that is a much better idea

terenzreignz (terenzreignz):

you and your formulas XD And I thought I was the one with that tendency ^_^

OpenStudy (anonymous):

\[\sum_{k=1}^n(2k)^2=4\sum_{k=1}^nk^2\]

terenzreignz (terenzreignz):

^And THAT is much neater ^_^

OpenStudy (anonymous):

i was getting there, just a little slower than necessary

OpenStudy (anonymous):

@terenzreignz great approach nw taking common 2^2 from each term we can say the expression is =4m

terenzreignz (terenzreignz):

Yup. Just factor out the 4, and you'll have the m left :D

OpenStudy (anonymous):

u too helped me a lot bro @satellite73

OpenStudy (anonymous):

yup i have got it nw bro thnq very much @terenzreignz

terenzreignz (terenzreignz):

That's all I need to know ^_^

OpenStudy (anonymous):

@satellite73 thnx to u too bro